OpenStudy (anonymous):

Need help Solving 6t - 2 = 0 This is how I laid it out: 6t -2 +2 = 0 +2 6t = 2 6t/6 = 2/6 2/6 = .33333333 t=.333333333 This cant be right - What am I doing wrong?

OpenStudy (anonymous):

y do u follow such tough methods?

OpenStudy (noseboy908):

no, you're right, just leave it as a fraction

OpenStudy (anonymous):

jst bring 2 to the opposite side bcoz u have to find the value of t.so it must be free from all coefficients 6t-2=0 6t=2 t=2/6 t=1/3

OpenStudy (anonymous):

6t-2=0 6t=2 t=2/6 t=1/3

OpenStudy (anonymous):

which equals to 0.333333...

OpenStudy (anonymous):

Ahhh! okay! Got it! I'll change to a fraction of 1/3. Thanks Everyone!

OpenStudy (anonymous):

u r welcome! :-)

OpenStudy (anonymous):

Just getting back in the saddle again after 20 years. So its baby steps for me trying to get a handle on it again! I appreciate everyone's help!